what is the final velocity of a 150 kg motorcycle driven by a 50.0 kg rider, accelerated from rest for 11 s by an unbalanced force of 800.0 N South
1
Expert's answer
2019-11-05T10:38:52-0500

The acceleration of the motorcycle


"a=\frac{F}{m}=\frac{800\:\rm N}{200\:\rm kg}=4\:\rm m/s^2"

The final velocity


"v_f=v_i+at=0+4\rm m/s^2\times 11\: s=44\: m/s"